# CoVR: Learning Composed Video Retrieval from Web Video Captions

### v3

@@ 4.2 Composed Video Retrieval Results
CompoDiff (Gu ‘23) ST18M 19.37 53.81 72.02 90.85
Pic2Word (Saito ‘23) CC3M 23.90 51.70 65.30 87.80 24.70 43.70
CASE (Levy ‘23) LaSCo 30.89 60.75 73.88 92.84 - -
CASE (Levy ‘23) LaSCo+COCO 35.40 65.78 78.53 94.63 - -
CoVR-BLIP - 19.95 41.98 52.31 73.54 16.00 32.77
CoVR-BLIP (Ours) WebVid-CoVR 39.28 68.22 78.94 94.65 27.70 44.63
@@
Table 3: State-of-the-art comparison on the CIRR test set and FashionIQ validation set: Our model benefits from training on WebVid-CoVR in both the zero-shot and finetuning setting. For full metrics, see Section C.8 of the Appendix. $\dagger$ denotes results reported by (Liu et al. 2021). For methods that pretrain specifically for composed retrieval, we indicate their pretraining data. CC3M denotes Conceptual Captions 3M (Sharma et al. 2018), ST18M denotes  (Gu et al. 2023), and COCO denotes (Lin et al. 2014).
We make several conclusions. (i) Combining both visual and text modalities yields better performance than the models with individual modalities. This result highlights that our new CoVR benchmark requires paying attention to both modalities. (ii) Visual-only outperforms text-only suggesting that the video pairs automatically mined through their caption similarity indeed exhibits visual similarity, and that the image captures the target video better than the modification text. (iii) Finetuning on WebVid-CoVR obtains substantial improvements over using pretrained and frozen embeddings. (iv) When finetuning, fusion with BLIP cross-attention (CA) performs better than the MLP fusion. (v) Results with the BLIP backbone are marginally higher than those with CLIP. (vi) Using $N=15$ target video frames further boosts the performance over using only the middle frame. We further analyse the effect of the number of frames in Section C.7 of the Appendix.
4.3 Transfer Learning to CoIR
While our focus is video retrieval, we also experiment with transferring our CoVR models to image retrieval tasks on standard CoIR benchmarks. We define zero-shot CoIR as not using any manually annotated CoIR triplet for training. We perform zero-shot CoIR by directly applying our model trained on our automatically generated WebVid-CoVR dataset to CoIR tasks and also explore finetuning our model on the training set of the downstream benchmark.
Table 3 report results on CIRR and Fashion-IQ datasets. These results show that our model highly benefits from training on WebVid-CoVR, especially in the zero-shot setting, on both datasets. In addition, our model achieves state-of-the-art performance in both zero-shot and finetuning settings on the CIRR and FashionIQ datasets.
@@ 4.4 Ablation Studies
We now ablate the importance of several key aspects of our method by training only on WebVid-CoVR.
@@
Importance of data scale. In Table 4, we evaluate the effect of the number of video-caption pairs used as a seed for our triplet generation pipeline. We construct subsets of videos such that larger ones include smaller ones, and only keep triplets that contain the sampled videos for training. We find that results steadily increase when using more videos, demonstrating that our method largely benefits from scaling the size of the seed dataset of video-captions. We also observe the importance of the filtering techniques described in Section 3.1, as the model trained on unfiltered data underperforms.
Initial Generated W-C-T CIRR FashionIQ
#videos #triplets Fil. MeanR MeanR MeanR
0 - - 36.80 46.95 24.39
200k 11k 62.20 67.01 37.04
500k 66k 76.12 69.92 36.98
1M 269k 77.89 70.03 36.47
2.5M 1.6M 79.40 70.27 36.17
2.5M 3.6M 78.92 69.25 32.31
@@
Table 4: Data size: We measure the importance of the number of videos used for data generation and of filtering (Fil. column) the generated data, by evaluating on WebVid-CoVR-Test (W-C-T), CIRR, and FashionIQ. All models are trained for the same number of iterations on the generated data.
@@
Modification text generation. We use a large language model finetuned for modification text generation (MTG-LLM) as explained in Section 3.1. We here compare this solution to a simple rule-based baseline that uses several templates to generate the modification text given the two captions that differ by one word. Specifically, the modification text is based on the two different words from the captions. We generate templates that use these words and choose one at random during training. These templates include variations such as ‘Remove $\mathtt{txt\_diff_{1}}$” and "Change $\mathtt{txt\_diff_{1}}$ for $\mathtt{txt\_diff_{2}}$. A full list of all the templates can be seen in Section B.3 of the Appendix. Additionally, we investigate the possibility of paraphrasing the rule-based modification texts using GPT-3.5-turbo from OpenAI (Brown et al. 2020) as a source of augmentation, by prompting “Paraphrase the following sentence: {Rule-base modification text}. In preliminary analysis, we qualitatively observed that LLaMA (Touvron et al. 2023) and LLaMA 2 (Touvron et al. 2023) alternatives were overly verbose when used for paraphrasing; however, GPT-3.5 outputs were satisfactory.
@@
In Table 5, we show that our MTG-LLM generates better modification texts than the rule-based baseline, by evaluating the results of the model trained on the generated data. Paraphrasing the rule-based examples significantly boosts the performance (from 41 to 52 R@1), while still being worse than our MTG-LLM, especially on the CIRR benchmark. Note that the paraphrasing comes with the cost of running an expensive LLM (\$43 cost for this experiment for 1 paraphrasing per modification text on the entire dataset). On the other hand, our MTG-LLM finetuning only requires 715 text examples. Qualitative examples comparing MTG-LLM and rule-based are provided in Table A.11 of the Appendix.
WebVid-CoVR-T. CIRR
Model R@1 MeanR R@1 MeanR
Rule-based 41.39 70.36 16.31 47.84
Rule-based paraph. 52.39 78.75 34.53 66.54
MTG-LLM 53.13 79.40 39.28 70.27
Table 5: Modification text generation: We compare our MTG-LLM (LLaMA 7B parameters) against both a rule-based MTG baseline and a paraphrased rule-based MTG baseline (using GPT-3.5-turbo from OpenAI). We observe important gains in the downstream performance of the model trained on the generated data.
WebVid-CoVR-T. CIRR
Iteration HN-NCE R@1 MeanR R@1 MeanR
Triplets 49.80 78.15 35.66 67.15
Videos 49.02 76.62 35.93 67.72
Videos 53.13 79.40 39.28 70.27
@@
Table 6: Training strategies: Iterating on batches of distinct target videos (instead of triplets) and up-sampling hard negatives (HN-NCE (Radenovic et al. 2023)) both benefit the CoVR/CoIR performance.
@@
Training strategies. In Table 6, we first show the benefit on WebVid-CoVR of training by iterating on target videos instead of CoVR triplets. This is to avoid having the same target video appearing multiple times in a training batch, hence increasing the number of correct negatives that are used in the contrastive loss. Furthermore, up-sampling hard negatives adopting the HN-NCE loss formulation from (Radenovic et al. 2023) also slightly benefits the performance.
@@ 5 Conclusions and Limitations
Our model constitutes a generic multi-modal search tool, but is not intended for a specific application. While there are helpful use cases such as online shopping, traveling, and personal development (i.e., how-to), there may be potential privacy and harmful risks when training the model on datasets with inappropriate content. The risks include surveillance applications such as searching for a specific person, and looking up violent and graphic videos. For our WebVid-CoVR data release, we refer to Section A for further analysis about removal of inappropriate content. We note that our dataset users must also adhere to the terms of use stipulated by WebVid (Bain et al. 2021).
@@ Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014223 made by GENCI. The authors would like to acknowledge the research gift from Google, the ANR project CorVis ANR-21-CE23-0003-01, Antoine Yang’s Google PhD fellowship, and thank Mathis Petrovich, Nicolas Dufour, Charles Raude, and Andrea Blazquez for their helpful feedback.
@@
This document provides dataset statistics (Section A), implementation details (Section B), additional experiments (Section C), and qualitative examples (Section D). We also provide the code and dataset together with a datasheet, and an illustrative video on our project page at imagine.enpc.fr/~ventural/covr.
Appendix A Dataset statistics and analysis
@@
We conduct a sentiment analysis on the modification texts using the TextBlob library (Loria 2013) to identify instances of negative sentiment. We find that less than 0.5% of the dataset (about 2k instances) exhibits negative sentiment. Upon manual review, we identify false positives in this categorization, including examples such as “make it an evil pumpkin” or “Change him into a frustrated businessman”. The instances detected as negative sentiment are reviewed and 260 of them are removed from the dataset. We ensure that the dataset does not include any videos marked for mature content, by checking the metadata of WebVid (Bain et al. 2021) provided by (Lab 2023). Finally, using the better-profanity library (Thanh 2018), we identify approximately 2k video captions that are marked for profanity. Upon manual inspection, we find that there were a large number of videos displaying computer-generated visuals with those words. We also notice false positives (e.g., misinterpretation due to context), such as the animal cock being incorrectly identified as profanity. The videos detected to contain profanity in their captions are reviewed and excluded from the dataset.
@@
Video categories. We plot the distribution of video categories in Figure A.4. These categories are found using the WebVid metadata provided by (Lab 2023). We find 50% of WebVid-CoVR videos in this metadata collection. Note more than one category can be associated with a single video (e.g., Nature and Animals/Wildlife for a video of a fish in the ocean).
@@
Figure A.4: Distribution of video categories: We plot the distribution of categories for videos in WebVid-CoVR, as provided by (Lab 2023) as WebVid metadata. Note that 50% of our WebVid-CoVR videos are present in this metadata collection. Looking at the distribution, we observe that around 40% and 20% of WebVid-CoVR are videos of Nature and People, respectively.
Figure A.5: Distribution of parts of speech in modification texts: Distribution of nouns, verbs, pronouns, adjectives, and adverbs in the modification text using part-of-speech (POS) tagging. On average, there are more than one noun and one verb per modification text.
@@
Appendix B Implementation details

### v4

@@ 4.4 State-of-the-art comparison on CoIR benchmarks
CompoDiff [6] 2023 ST18M+LAION2B 26.71 55.14 74.52 92.01 64.54 82.39 91.81 37.78 49.10 41.31 55.17 44.26 56.41 39.02 51.71
Pic2Word [21] 2023 CC3M 23.90 51.70 65.30 87.80 - - - 20.00 40.20 26.20 43.60 27.90 47.40 24.70 43.70
SEARLE-XL-OTI [4] 2023 CC3M 24.87 52.31 66.29 88.58 53.80 74.31 86.94 21.57 44.47 30.37 47.49 30.90 51.76 27.61 47.90
SEARLE-XL [4] 2023 CC3M 24.24 52.48 66.29 88.84 53.76 75.01 88.19 26.89 45.58 20.48 43.13 29.32 49.97 25.56 46.23
GRB [23] 2023 - 24.19 52.07 65.48 85.28 60.17 79.13 89.34 24.14 45.56 34.54 55.15 33.55 53.60 30.74 51.44
GRB+LCR [23] 2023 - 30.92 56.99 68.58 85.28 66.67 78.68 82.60 - - - - - - - -
TFCIR [23] 2023 - 32.82 61.13 71.76 85.28 66.63 78.58 82.68 24.84 45.56 35.38 55.15 33.10 53.60 31.11 51.44
CASE [7] 2024 LaSCo 30.89 60.75 73.88 92.84 60.17 80.17 90.41 - - - - - - - -
CASE [7] 2024 LaSCo+COCO 35.40 65.78 78.53 94.63 64.29 82.66 91.61 - - - - - - - -
CIReVL [25] 2024 - 34.65 64.29 75.06 91.66 67.95 84.87 93.21 27.07 49.53 33.71 51.42 35.80 56.14 32.19 52.36
LinCIR (ViT-L) [24] 2024 - 25.04 53.25 66.68 - 57.11 77.37 88.89 20.92 42.44 29.10 46.81 28.81 50.18 26.28 46.49
LinCIR (ViT-G) [24] 2024 - 35.25 64.72 76.05 - 63.35 82.22 91.98 38.08 60.88 46.76 65.11 50.48 71.09 45.11 65.69
CoVR-BLIP [58] 2024 WebVid-CoVR 38.48 66.70 77.25 91.47 69.28 83.76 91.11 21.95 39.05 30.37 46.12 30.78 48.73 27.70 44.63
CoVR-BLIP-2 2024 CC-CoIR 43.35 73.78 83.66 96.07 75.25 88.89 95.23 33.17 55.13 38.81 58.24 37.48 58.49 36.49 57.29
CoVR-BLIP-2 2024 WV-CC-CoVIR 43.74 73.61 83.95 96.10 72.84 87.52 94.39 34.26 56.22 41.22 59.32 38.96 59.77 38.15 58.44
@@
TABLE V: State-of-the-art comparison on the CIRCO test set: Our model benefits from training on WV-CC-CoVIR in the zero-shot setting. For methods that pretrain specifically for composed retrieval, we indicate their pretraining data. CC3M denotes Conceptual Captions 3M [13] and ST18M denotes [6].
Pretraining mAP@K
Method Year Data K=5 K=10 K=25 K=50
Random - 00.00 00.00 00.00 00.00
SEARLE-XL [4] 2023 CC3M 11.68 12.73 14.33 15.12
CompoDiff (ViT-G) [6] 2023 ST18M 15.33 17.71 19.45 21.01
GRB+LCR [23] 2023 - 25.38 26.93 29.82 30.74
TFCIR [23] 2023 - 26.52 28.25 31.23 31.99
CIReVL [25] 2024 - 26.77 27.59 29.96 31.03
LinCIR (ViT-G) [24] 2024 - 19.71 21.01 23.13 24.18
CoVR-BLIP [58] 2024 WebVid-CoVR 21.43 22.33 24.47 25.48
CoVR-BLIP-2 2024 CC-CoIR 23.50 24.66 27.05 28.04
CoVR-BLIP-2 2024 WV-CC-CoVIR 28.29 29.55 32.18 33.26
TABLE VI: Adding the composed caption retrieval loss: We compare models with and without using our additional caption retrieval loss, on both CoVR-BLIP [58] and CoVR-BLIP-2. We observe a significant improvement on CoIR datasets when training on the WebVid-CoVR with this additional loss.
Caption WebVid-CoVR-T. CIRR FashionIQ CIRCO
Method ret. loss R@1 R@1 R@10 mAP@5
CoVR-BLIP 56.81 37.18 23.20 22.86
CoVR-BLIP 55.95 40.46 33.90 25.71
CoVR-BLIP-2 60.09 34.43 27.78 25.06
CoVR-BLIP-2 59.82 41.42 36.81 27.84
We report results in Table IV on CIRR and Fashion-IQ datasets, and in Table V for the newer CIRCO dataset. In Table IV, we report both finetuning (top block), and zero-shot (bottom block) settings. In Table V, we only report the zero-shot setting since CIRCO does not have a training set to finetune on. These results show that our model highly benefits from training on WV-CC-CoVIR, especially in the zero-shot setting. Remarkably, our model attains state-of-the-art performance on CIRR and CIRCO, exhibiting a 5% improvement in CIRR (43.74 vs 38.48 R@1) and a 2% in CIRCO (28.29 vs 26.77 mAP@5) compared to the nearest competitor’s results. As for the FashionIQ dataset, we outperform all other methods but CompoDiff [6] and LinCIR [24] with the ViT-G backbone. With CompoDiff, the results are mixed depending on the metric (38.15 vs 39.02 R@1 and 58.44 vs 51.71 R@10). For LinCIR, if we compare with the ViT-L backbone, more similar to ours, our methods obtains better results (38.15 vs 26.28). However, CompoDiff/LinCIR perform poorly on the other two datasets CIRR (43.74 vs 26.71/35.25) and CIRCO (28.29 vs 15.33/19.71). Our CoVR-BLIP-2 remains therefore the overall best zero-shot model when evaluating across three datasets. In addition to this strong zero-shot performance, our model reaches state-of-the-art performance when finetuned on both CIRR and FashionIQ benchmarks (top block of Table IV). For each setting, we also provide results with pretraining only on the CC-CoIR subset, as opposed to the full WV-CC-CoVIR, and observe similar performance on CIRR and FashionIQ, but lower on the challenging CIRCO dataset.
4.5 Comparison with synthetic training images
To further compare against other approaches that propose automatic triplets for training (i.e., by generating synthetic target images [11, 6]), we train our CoVR-BLIP-2 model on their data in a controlled experiment. We summarize the results in Table VII. In the top block, we show numbers directly taken from the CompoDiff [6] work, comparing their method trained on various synthetic datasets: InstructPix2Pix dataset of 1-million triplets [11], a 1M subset of their SynthTriplets [6], and the full 18M version. Note that we use the same metrics as in their paper (e.g., average of R@1, Rs@1 for CIRR) and show the results with their ViT-L model. In the bottom block, we train our model on the 1M synthetic dataset versions, as well as a 1M subset of our real visual data from WV-CC-CoVIR. When comparing the two models on the same training data, we observe similar performances (e.g., 28.32 vs 28.44 on CIRR), suggesting that the main difference comes from training data, rather than the model. Training on our WV-CC-CoVIR with 1M triplets outperform SynthTriplets with both 1M and 18M versions, highlighting the importance of real training images.
TABLE VII: Synthetic vs real training images: We compare the CoIR performance of our proposed method (CoVR-BLIP-2) and dataset (WV-CC-CoVIR) against the CompoDiff [6] method and their proposed dataset (SynthTriplets). The results demonstrate that our training data achieves better performance with 1M triplets compared to 1M or even 18M triplets of the SynthTriplets dataset, containing synthetically generated target images. IP2P denotes InstructPix2Pix, 1M public synthetic triplets by [11].
Pretraining Data WebVid-CoVR-T CIRR FashionIQ
Model Data Size R@1 Avg(R@1, Rs@1) Avg(R@10, R@50)
CompoDiff [6] IP2P [11] 1M - 27.42 27.24
SynthTriplets [6] 1M - 28.32 31.91
SynthTriplets [6] 18M - 37.83 42.33
CoVR-BLIP-2 IP2P [11] 1M 16.55 34.72 15.42
SynthTriplets [6] 1M 45.42 28.44 33.23
WV-CC-CoVIR 1M 55.87 58.79 48.14
TABLE VIII: Modification text generation: We compare our finetuned model MTG-LLM (LLaMA 7B parameters) against (a) a rule-based MTG baseline, (b) a paraphrased rule-based MTG baseline (using GPT-3.5-turbo from OpenAI), and (c) simply prompting the frozen LLaMA LLM. We observe important gains in the downstream performance of the model trained on the generated data.
WebVid-CoVR-Test CIRR
Model R@1 R@5 R@10 R@50 R@1 R@5 R@10 R@50
Rule-based 39.08 67.33 78.13 93.82 12.02 34.75 47.06 75.35
Rule-based & GPT-paraphrased 57.94 81.85 89.79 97.93 32.89 61.98 73.04 90.34
Prompting LLaMA 56.46 82.08 89.32 97.85 34.27 64.29 75.76 91.61
MTG-LLM 59.82 83.84 91.28 98.24 41.42 72.58 82.55 96.31
@@ 4.6 Ablation studies
We now ablate the importance of several key aspects of our method by focusing primarily on experiments trained in WebVid-CoVR, and also examine the impact of different datasets on model performance.
The additional composed caption retrieval loss. As explained in Section 3.3, we integrate a caption retrieval loss term as additional supervision. For both methods (CoVR-BLIP [58] and CoVR-BLIP-2), this led to a significant improvement in the CoIR performance and a slight decrease on WebVid-CoVR on which it was originally trained on, see Table VI (34 vs 41 R@1 on CIRR and 27 vs 36 R@1 on FashionIQ for instance).
@@
Importance of data scale. In Table IX, we evaluate the effect of the number of video-caption pairs used as a seed for our triplet generation pipeline. We construct subsets of videos such that larger ones include smaller ones, and only keep triplets that contain the sampled videos for training. We find that results steadily increase when using more videos, demonstrating that our method largely benefits from scaling the size of the seed dataset of video-captions. We also observe the importance of the filtering techniques described in Section 3.1, as the model trained on unfiltered data underperforms.
@@
TABLE IX: Data size: We measure the importance of the number of videos used for data generation and of filtering the generated data, by evaluating on WebVid-CoVR-Test, CIRR, and FashionIQ. All models are trained for the same number of iterations on the generated data.
Initial Generated WebVid-CoVR-Test CIRR FashionIQ CIRCO
#videos #triplets Filtering R@1 MeanR R@1 MeanR R@10 R@50 mAP@5
0 - - 16.55 36.15 18.60 47.82 09.75 21.09 04.83
200k 11k 51.53 77.92 40.72 70.82 34.51 55.82 24.00
500k 66k 55.13 80.59 40.22 70.67 36.04 57.10 25.86
1M 269k 57.32 82.13 40.55 71.08 36.76 57.13 26.92
2.5M 1.6M 59.82 83.30 41.42 73.22 36.81 56.70 27.84
2.5M 3.6M 58.65 82.57 40.84 71.11 37.19 57.19 29.11
@@
Modification text generation. We use a large language model finetuned for modification text generation (MTG-LLM) as explained in Section 3.1. We here compare this solution to prompting it without any training and to a simple rule-based baseline that uses several templates to generate the modification text given the two captions that differ by one word. For prompting, we prepend few-shot examples of pairs of captions and desired generated texts, before adding the two captions in question. Please check Section C.5 of the Appendix for the full prompt. Table VIII shows that finetuning the MTG-LLM for generating the training data is much more effective than prompting it without finetuning, as measured by CoVR performance on WebVid-CoVR-Test and CoIR performance on CIRR. For the rule-based experiment, the modification text is based on the two different words from the captions. We generate templates that use these words and choose one at random during training. These templates include variations such as ‘Remove $\mathtt{txt\_diff_{1}}$” and "Change $\mathtt{txt\_diff_{1}}$ for $\mathtt{txt\_diff_{2}}$. A full list of all the templates can be seen in Section C.3 of the Appendix. Additionally, we investigate the possibility of paraphrasing the rule-based modification texts using GPT-3.5-turbo from OpenAI [71] as a source of augmentation, by prompting “Paraphrase the following sentence: {Rule-base modification text}. In preliminary analysis, we qualitatively observed that LLaMA [10] and LLaMA 2 [72] alternatives were overly verbose when used for paraphrasing; however, GPT-3.5 outputs were satisfactory.
@@
In Table VIII, we show that our MTG-LLM generates better modification texts than the rule-based baseline, by evaluating the results of the model trained on the generated data. Paraphrasing the rule-based examples significantly boosts the performance (from 39 to 57.9 R@1), while still being worse than our MTG-LLM, especially on the CIRR benchmark. Note that the paraphrasing comes with the cost of running an expensive LLM (\$43 cost for this experiment for 1 paraphrasing per modification text on the entire dataset). On the other hand, our MTG-LLM finetuning only requires 715 text examples. Qualitative examples comparing MTG-LLM and rule-based are provided in Table A.11 of the Appendix.
@@
TABLE X: Training strategies: Iterating on batches of distinct target videos (instead of triplets) and up-sampling hard negatives both benefit the CoVR/CoIR performance.
HN-NCE WebVid-CoVR-Test CIRR
Iteration [15] R@1 R@5 R@10 R@50 R@1 R@5 R@10 R@50
Triplets 53.79 81.69 88.97 97.89 37.49 66.63 77.11 91.28
Videos 55.24 81.34 89.48 98.24 38.70 68.84 78.41 92.72
Videos 59.82 83.84 91.28 98.24 41.42 72.58 82.55 96.31
@@
Training strategies. In Table X, we first show the benefit on WebVid-CoVR of training by iterating on target videos instead of CoVR triplets. This is to avoid having the same target video appearing multiple times in a training batch, hence increasing the number of correct negatives that are used in the contrastive loss. Furthermore, up-sampling hard negatives adopting the HN-NCE loss formulation from [15] also slightly benefits the performance.
4.7 Qualitative analysis
In this section, we provide qualitative examples of our WebVid-CoVR and CC-CoIR triplets. Figure 5, shows examples of triplets generated using our automatic dataset creation. These examples demonstrate the effectiveness of our approach in generating coherent modification texts for paired videos. For more examples, we refer to Section E.5 of the Appendix.
Figure 5: Examples of generated triplets: We illustrate triplet samples generated using our automatic dataset creation methodology (left: WebVid-CoVR, right: CC-CoIR). Each sample consists of two videos/images with their corresponding captions (at the bottom of each video/image) and the generated modification text using our MTG-LLM (in purple).
@@ 5 Conclusions and Limitations
Our model constitutes a generic multi-modal search tool, but is not intended for a specific application. While there are helpful use cases such as online shopping, traveling, and personal development (i.e., how-to), there may be potential privacy and harmful risks when training the model on datasets with inappropriate content. The risks include surveillance applications such as searching for a specific person, and looking up violent and graphic videos. For our WebVid-CoVR data release, we refer to Section A for further analysis about removal of inappropriate content. We note that our dataset users must also adhere to the terms of use stipulated by WebVid [12].
@@ Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014223 made by GENCI. The authors would like to acknowledge the research gift from Google, the ANR project CorVis ANR-21-CE23-0003-01, Antoine Yang's Google PhD fellowship, and thank Mathis Petrovich, Nicolas Dufour, Charles Raude, and Andrea Blazquez for their helpful feedback.
@@
[…]
This document provides WebVid-CoVR dataset statistics (Section A), CC-CoIR dataset statistics (Section B), implementation details (Section C), additional experiments (Section D), and additional qualitative examples (Section E). We also provide the code and dataset together with a datasheet, and an illustrative video on our project page at imagine.enpc.fr/~ventural/covr.
A WebVid-CoVR dataset statistics and analysis
@@
We conduct a sentiment analysis on the modification texts using the TextBlob library [73] to identify instances of negative sentiment. We find that less than 0.5% of the dataset (about 2k instances) exhibits negative sentiment. Upon manual review, we identify false positives in this categorization, including examples such as “make it an evil pumpkin” or “Change him into a frustrated businessman”. The instances detected as negative sentiment are reviewed and 260 of them are removed from the dataset. We ensure that the dataset does not include any videos marked for mature content, by checking the metadata of WebVid [12] provided by [74]. Finally, using the better-profanity library [75], we identify approximately 2k video captions that are marked for profanity. Upon manual inspection, we find that there were a large number of videos displaying computer-generated visuals with those words. We also notice false positives (e.g., misinterpretation due to context), such as the animal cock being incorrectly identified as profanity. The videos detected to contain profanity in their captions are reviewed and excluded from the dataset.
@@
Video categories. We plot the distribution of video categories in Figure A.4. These categories are found using the WebVid metadata provided by [74]. We find 50% of WebVid-CoVR videos in this metadata collection. Note more than one category can be associated with a single video (e.g., Nature and Animals/Wildlife for a video of a fish in the ocean).
@@
Figure A.4: Distribution of video categories: We plot the distribution of categories for videos in WebVid-CoVR, as provided by [74] as WebVid metadata. Note that 50% of our WebVid-CoVR videos are present in this metadata collection. Looking at the distribution, we observe that around 40% and 20% of WebVid-CoVR are videos of Nature and People, respectively.
Figure A.5: Distribution of parts of speech in modification texts: Number of nouns, verbs, pronouns, adjectives, and adverbs in the modification text using part-of-speech (POS) tagging. On average, there are more than one noun and one verb per modification text.
@@
WebVid-CoVR dataset overlap with zero-shot CoIR evaluation datasets. To contextualize the zero-shot performance, we analyze the potential overlap between our WebVid-CoVR dataset and the three CoIR datasets. We compute the CLIP [2] embeddings for all target videos in the CC-CoIR training set and calculate their similarity with the target images in the test sets of each CoIR dataset. We then evaluate the overlap by setting similarity thresholds at 0.7, 0.8, and 0.9. We define overlap as occurring when at least one sample in the test set has a similarity score above the threshold. We note that for simplicity, we only consider target images.
The results, summarized in Table A.1, show that no target videos from the WebVid-CoVR dataset exhibit a similarity score higher than 0.9 with any of the CoIR datasets, indicating zero overlap at the highest threshold. However, as the threshold decreases, the overlap increases, specially for CIRCO, which shows a 12.6% overlap at a 0.8 threshold.
TABLE A.1: Overlap between WebVid-CoVR training data and zero-shot CoIR benchmarks: We present the percentage overlap between target videos in our WebVid-CoVR training dataset and the target images in the test sets of three CoIR datasets at different CLIP cosine similarity thresholds (0.7, 0.8, and 0.9). Overlap is defined as the presence of at least one target image in the test set with a similarity score above the specified threshold. The results indicate no overlap at the highest threshold (0.9). Note that this analysis focuses on the overlap of target images, not triplets.
FashionIQ
Threshold CIRR Dress Shirt Toptee CIRCO
0.7 45.4% 13.6% 13.8% 10.9% 79.0%
0.8 01.6% 00.3% 00.0% 00.0% 12.6%
0.9 00.0% 00.0% 00.0% 00.0% 00.0%
B CC-CoIR dataset statistics
In this section, we provide analysis on our CC-CoIR. We start with 3.3M caption-image pairs, with 2M distinct captions. As explained in Section 3.1 of the main paper, we mine paired images by searching for captions that differ by a single word, excluding punctuation marks. This process allows us to identify a vast pool of 1.2M distinct caption pairs with 281k distinct captions, resulting in 3.3M triplets after filtering.
CC-CoIR image categories. Figure A.8 illustrates the distribution of the top 20 categories derived from our CC-CoIR dataset. We find 67% of images within our CC-CoIR dataset possess one or more associated categories. It’s worth mentioning that a single video may be associated with multiple categories simultaneously. For instance, a video featuring a singer may be categorized under both "Singer" and "Music Artist". It is important to note that while only the top 20 are displayed, the complete dataset encompasses over 10,000 distinct categories. This highlights the wide variety of visual content present within our collection.
Figure A.8: Distribution of image categories (Top 20) in CC-CoIR: We plot the distribution of categories for images in CC-CoIR, as provided by [13]. Note that 67% of our CC-CoIR images have one or more categories in this metadata collection. Looking at the distribution, we observe that around 25% and 15% of WebVid-CoVR are videos of illustration and art, respectively.
Quantifying noise in CC-CoIR. Here, we attempt to quantify how well the modification text describes the transformation of the query image into the target image. While this task is challenging, one possible approach involves leveraging the InstructPix2Pix [11] model. We input the source image and corresponding modification text into the model to generate an image that reflects the intended transformation. The results, summarized in Figure A.9, reveal that the majority of target images have a cosine similarity “close” to that of the generated images. This suggests that the modification text generally aligns with the visual changes captured in the target images. However, a qualitative examination of the generated images indicates that they cannot always serve as ground truth, suggesting that these results should be interpreted with caution.
Figure A.9: Cosine similarity histogram between target image and generated image: We plot the distribution of cosine similarity scores between target images and images generated using InstructPix2Pix [11] based on the corresponding source image and modification text. The results suggest that the majority of target images exhibit a similarity to the generated images.
CC-CoIR dataset overlap with zero-shot CoIR evaluation datasets. As previously done for WebVid-CoVR, we now analyze the overlap between our CC-CoIR dataset and the CoIR datasets using CLIP [2] embeddings. The results, summarized in Table A.2, reveal that fewer than 2% of the target images in the CC-CoIR training set have a similarity score higher than 0.9 in any of the CoIR datasets. When the threshold is relaxed to 0.8, the overlap increases, with CIRCO showing 20.7% and CIRR showing 5.2% of samples with a similarity score above the threshold.
TABLE A.2: Overlap between CC-CoIR training data and zero-shot CoIR benchmarks: We repeat the overlap analysis with CC-CoIR (as is similarly done for WebVid-CoVR in Table A.1). The results indicate minimal overlap at the highest threshold (0.9), with fewer than 2% similarity across all datasets, suggesting limited direct overlap.
FashionIQ
Threshold CIRR Dress Shirt Toptee CIRCO
0.7 34.0% 06.9% 10.8% 09.6% 59.1%
0.8 05.2% 00.2% 00.3% 00.2% 20.7%
0.9 00.2% 00.0% 00.0% 00.0% 01.6%
C Implementation details